Presentation is loading. Please wait.

Presentation is loading. Please wait.

1 Virgo commissioning progress Post VSR1 plans E. Tournefier STAC meeting June 26 th,2007.

Similar presentations

Presentation on theme: "1 Virgo commissioning progress Post VSR1 plans E. Tournefier STAC meeting June 26 th,2007."— Presentation transcript:

1 1 Virgo commissioning progress Post VSR1 plans E. Tournefier STAC meeting June 26 th,2007

2 2 carrierSidebands 15 minutes Status at last STAC Thermal effects => fine tuning of parameters during lock acquisition + slow lock acquisition ITF stably locked (Weekend Science Runs) but Controls needed to be improved (longitudinal, angular, suspension): => stability, noise re-introduction, stationarity, duty cycle Control noises had to be reduced and environmental noises to be understood WSR5 (Nov 2006)

3 3 Status now Thermal effects still with us (waiting for thermal compensation) => still need fine tuning of parameters during lock acquisition + slow lock acquisition ITF control well improved: –reliable and more stable controls (long locks), good stationarity, –control noise subtraction (no impact above 40 Hz), –duty cycle improved against environmental conditions Noise well reduced (both controls and environmental) but… Still noises to understand

4 4 Next steps – Diffused light  Install as soon as possible acoustic isolation  Detection lab  End benches  Perform diffused light mitigation  Detection lab  End benches  Understand better a solution for the Brewster noise  Replacement with a bigger one  Remove it ?  Understand possible clipping/spurious beam in vacuum  Modeling of all the beams  Mitigation – conservative approach OK Done Shall we? OK (to be continued) Last STAC

5 5 Next steps – Control noise  Frequency noise improvements  Install Better SSFS (common mode servo ) electronics (under design)  Understand sensing noise (diffused light?)  Longitudinal control noise reduction  More aggressive filtering  Alpha technique frequency dependent  Understand sensing noise (diffused light?) and couplings in the auxiliary d.o.f. error signal (thermal lensing?)  Angular degrees of freedom  Mirror/beam centering – coils balancing ongoing  More aggressive filtering on going OK started OK OK (improved error signal) Last STAC

6 6 Next steps – Control improvements  Increase gain of the automatic alignment loops  Improve suspension control  Global inverted pendulum control  Vertical inertial damping OK OK (and more) Last STAC

7 7 Activities since last STAC December 2006 – April 2007: 5 more WSRs –Improvements of controls (longitudinal, angular, suspensions) –Environmental noise (diffused light, magnetic) investigations and reduction April 6 th – 18 th electrical shutdown: Infrastructure works: –Electrical works, Computing (UPS, civil work, re-organisation,..) Detector improvements -Acoustic enclosure around external detection bench -Replacement of Brewster window with a larger one Restart went ok: ITF relocked within ~2days April 18 th –May 18 th : –Acoustic enclosures at end benches –Scattered light investigations (Brewster) + other environmental noises –Improvements of controls (longitudinal, angular, suspensions) –Change of naming convention (V1:channel_name)for easier data exchange with LSC May 18 th: start VSR1 –Continue noise investigations + small control improvements

8 8 Interferometer controls Thermal side-effects Longitudinal and angular controls Mirror suspension control

9 9 Lock acquisition: thermal transient Fine tuning needed for lock acquisition: –demodulation phases, engagement of loops, gains, offsets on error signals,… Plus: Variations in the thermal transient: seem related to activities in laser lab but not understood  Some lengthy relocks

10 10 Thermal effects on sidebands: first studies Image of the sidebands (first step of phase camera): scanning system on B5 beam: - the beam is scanned on a pin-hole + photodiode - demodulated signal is used to reconstruct the sidebands shape  Clear defocusing effect  Need thermal compensation system + complete phase camera (see Michele’s talk)

11 11 Mirror longitudinal controls: noise reduction PRCL control noise => sensed by B5 => MICH control noise => sensed by dark fringe (B1) 1/ Built a signal as stable AND less noisy than B2_3f (mixing B2 and B2_3f) 2/ online noise subtraction: - include frequency dependence  (f) - add a servo to tune the time dependance of the global gain   PRCL  MICH B2_ACpmix  (f)  (f) => MICH noise impact reduced by ~ 50 - with  =0 - with measured  (f)

12 12 Frequency stabilization and frequency noise New electronics designed and installed: –Better layout –Improved filter => frequency noise reduced at high frequencies + new functionality (switch between input signals) Alignment stability improved  coupling of frequency noise also well reduced

13 13 Mirror actuator noise  BS actuator noise reduced in May  The arm mirrors actuators need to be improved: to be done during VSR1 (will we find the Eddy currents noise?)  Next: new coil drivers + new DACs (see Michele’s talk) DAC DSP CoilDriver G EmphasisDeEmphasis

14 14 Longitudinal control noises: summary WSR5 (Nov 2006) VSR1 (now) MICH noise PRCL noise frequency noise

15 15 Mirror angular controls Control loops improved: –Add a second modulation (8 MHz) to control the common end motion –Optimized filters with more gain => robustness and stability Global control upgrade: more flexibility, noise injection,… Alignment signals DF and recycled powers 8MHz -> Drift control (10mHz) -> Drift control

16 16 Angular control noise Beam centering on the mirrors: –Now within ~ 1mm => reduced coupling –Now automated More aggressive filters  less noise reintroduction Angular noise budget  The angular noise is not limiting the current Virgo sensitivity This noise can be further reduced: -More power on B1p quadrant -Electronic noise reduction: new electronics (see Michele’s talk)  Should be compliant with Virgo design above 20-30 Hz B7 B8_q2 B8 WE WI NI NE BS

17 17 Suspension control: bottom stage Compensation of non-linear z ->  coupling: Large seismic activity  large correction sent to the marionette (zM)  y recoil of F7 (and payload)  Side effects on alignment Solution: compensate for z ->  y coupling Need a quadratic compensation: Dark fringe: alignment and total power - No compensation - With compensation  y=a*zM +b*zM 2 (payload: b=a/25) Correction signals zM  ycorr yy Under similar seismic activity:

18 18 Suspension control: top stage Old problem: environmental noise reintroduced by Lvdt (  seism) and Acc (wind) used for top stage control Several strategies developed: 1/ Panoply of error-signal blending depending on environmental conditions (‘onfly tuning’) 2/ Use the ITF longitudinal error signal instead (Global Inverted Pendulum Control) => z correction to the mirror well reduced GC (reconstructed z) LVDT VSR1: enabled only for NE-WE No GIPC With GIPC oldnew Other improvements: Vertical damping on long suspensions Local controls: optimized filters

19 19 Sensitivity versus wind and sea activity Illustration of the improvements of angular control and suspension control Empirical formula: Horizon = H 0 (1 - a x  seism – b x wind) WSR1 (Sept 2006): a= 0.37, b= 0.37VSR1 (May-Jun 2007):a= 0.07, b= 0.04  Sensitivity to bad weather conditions is well reduced  The lock can be kept in bad weather conditions - Measured horizon - Sea activity - Wind activity - Predicted horizon

20 20 Environmental noises magnetic noise scattered light - optical benches - Brewster windows

21 21 Magnetic noise Before WSR10 magnetic noise was limiting the sensitivity between 50 and 110 Hz Investigations: track the sources of magnetic field close to the input mirrors (more sensitive due to wrong polarity of the magnets) => found noisy power supplies => power supplies displaced Home made portable magnetometer Dark fringe Still some evidence for magnetic noise: under investigation

22 22 Scattered light Environmental noise acting on in air components + Scattered light from in air components = ------------------------------------------------------------- Phase noise in the interferometer (with the typical environmental-like noise structures) Laser Brewster windows End benches External bench Injection bench Detection suspended bench Actions: - reduce diffused light on benches - acoustic isolations - large Brewster windows

23 23 Scattered light at end benches Evidence for diffused light by the optics of the end benches (Jan 2007) Actions: 1/ use more rigid mounts for critical optics and dump secondary beams (Feb 2007) WSR7 WSR9: after 1/ Acoustic injections => acoustic noise still very close to the sensitivity floor 2/ install acoustic enclosures Tentative acoustic noise projection

24 24 Acoustic enclosures External detection bench: –Displace bench from the tower (~50cm) –Install acoustic panels (a room into the room) End benches: –Simpler: build a room around the benches Detection End buildings

25 25 Scattered light and acoustic isolation Tentative acoustic noise projection with acoustic enclosure Seismic noise on the optical bench - No acoustic enclosure - With acoustic enclosure After acoustic enclosure installation: Impact of acoustic noise well reduced above 100 Hz Tentative acoustic noise projection before

26 26 Brewster window + detection tower Brewster link Detection tower Tentative projection of ‘Brewster noise’ Observations: - Brewster+detection tower = area very sensitive to acoustic noise - Tentative noise projections: Brewster could be a limiting factor But: very difficult to make the difference between the Brewster and the detection tower  Damp all spurious beams inside the detection tower  Replace Brewster with a larger one (to fulfill: diameter > 5 x beam waist)

27 27 Detection tower scattered light Dump the secondary beams with black glass baffles: - Reflection by output port windows - Secondary beams (from second face of optics) on the suspended bench B1s B1 B1p B5 L3 /2 OMC M1 M5 L7

28 28 New Brewster window April shutdown: Detection Brewster replaced with a larger one: 16cm -> 19.6 cm Result: the noise is increased… After testing several hypothesis, found the culprit: aluminium baffle => Protect it from light the retro-diffused by the detection bench Detection

29 29 New Brewster window ‘patched’ After protecting the aluminium baffle: recovered the pre-shutdown sensitivity Additionnal small improvements: - ‘damp’ the Brewster link (weight+rubber) - smaller vacuum pump on SR tower  improve sensitivity from 100 to 200Hz The path of the noise in this area still has to be understood Old Brewster New Brewster ‘improved’ New Brewster

30 30 Sensitivity and noise budget

31 31 Sensitivity measurement: actuators calibration Found a frequency dependence of the actuators response (could be due to Eddy current effects)  “Improves” sensitivity by 20% at high frequencies 10%

32 32 VSR1 noise budget Magnetic noise: under investigation Not yet understood: hints for jitter/Pnoise from the injection system Environmental noise (laser lab) Actuator noise => shaping filters to be installed Laser power OMC matching

33 33 Environmental noises: some hints Below 30 Hz: beam jitter due to acoustic noise on laser bench ? some improvements obtained with better air conditioning flux Structures from 200 to 800Hz: input beam jitter, power noise?  Investigations ongoing during the run Input beam jitter Acoustic noise on laser bench Beam jitter vs acoustic noiseBeam jitter vs dark fringe dark fringe IMC reflection

34 34 Detector operation – Few words on VSR1

35 35 Detector monitoring/automation Detector monitoring, automation: the essential every day’s life tools, regularly upgraded Logbook: regular additions depending on user needs (more with the run!)

36 36 VSR1 : some statistics Good stability, long locks ( up to 59 hours) Gaussianity of the data is good, low trigger rate, low SNR NS-NS horizon (averaged) ~ 3.5 Mpc Science mode duty cycle ~ 85% Unlocks: –Technical: IMC fast unlock, Gc crash –Earthquakes, (bad weather) –Maintenance, commissioning breaks Horizon NS-NS (averaged value) 1 month

37 37 Commissioning activities during VSR1 Alignment drifts observed during long locks: related to the mis-centering of one end quadrant (used for BS control) => spoils error signal (AC)  Improvement: include quadrant asymmetry in the error signal for BS control Small drifts still there  Try to use another error signal (B1p?) Longer term:  Install galvanometer to keep centering in science mode  Reshuffle the end benches to have cleaner signals realignment Dark fringe Q81 asymmetry IMC fast unlocks - adjustement of IMC gain loop - repair electronics (loose connections) => rate smaller Suspension controls: some improvements /test of configurations / try to keep lock during earthquakes Environmental noise investigations: magnetic, acoustic, seismic

38 38 Long term observations: Etalon effect! Etalon effect: small FP cavity inside input mirrors (due to AR face)  effective reflectivity modulated with mirror thickness (temperature) … and so the finesse B7/B8 phd -FP transmitted power -Input mirror temperature Temperature variation FP Transmitted power No clear effect on sensitivity now … but to take into account for Virgo+   F/F = ± 3.5%

39 39 Post VSR1 plans Noises: Environmental noise investigations are going on during VSR1, simple actions will be done during the run –Magnetic noise close to the mirrors: identify sources and ‘remove’ them –Acoustic, seismic noise in the laser lab: find noise sources, reduce coupling to ITF –Acoustic noise at detection port: understand the path => remove Brewster? -Reduce the mirror actuator noise: shaping filters (to be done during VSR1) -Angular noise reduction (not urgent) Controls: –Rearrange the end benches: better error signals from quadrants –Install better centering system for quadrants (‘science mode’ compatible) –Improve the control of the short suspensions (now the limiting ones) –Revise the longitudinal control using the 8MHz modulation? (cleaner signals) Thermal effects: (see Michele’s talk) –Clean the input mirrors ? (if not too risky) –Install and commission the thermal compensation system when ready (Jan 2008?) + phase camera

40 40

41 41 Conclusion A lot of progress on the interferometer controls (still suffering from thermal effects): –Stability, robustness, duty cycle  ok for science data taking –Noise re-introduction  only a limitation below 40 Hz –Control scheme can still be improved: Quadrants / use 8MHz for longitudinal / short suspensions Good progress on the environmental noise: –Magnetic noise: well reduced  still under investigation –Scattered light: Benches: improvement of benches setup + acoustic enclosures  End benches and detection bench are safe now Brewster: larger one installed + investigations on noise path  To be better understood (remove Brewster?) –Remaining environmental noise under investigation (ISYS)  Actions to be defined when we have clearer ideas

42 42 Backup slides

43 43 Discussion ITF noise larger because of new Brewster Brewster link vibrating more than before? No. Sensitivity to Brewster vibrations larger than before? Yes Why? What is the mechanisms ? BEAM SPLITTER DETECTION 1) Scattering of spurious light coming from the ITF 2) Scattering of spurious light generated in the Brewster 3) Scattering of spurious light coming from the detection

44 44 With new larger Brewster new light visible on the detection bench - Light visible at ~10 cm from the axis of the main beam - This light is missing the first detection lens (L1, diameter 12 cm) - This light is diffused on a couple of supports on the detection bench New observations (III) B1s B1 B1p B5 L3 /2 OMC M1 M5 L7

45 45 Julien Marque (2004) Old prediction: spurious beams Bd3 Bd4 Bi1

46 46 Julien Marque (2004) Old prediction: spurious beams Bd3 and 4

47 47 With the old brewster One of the predicted spurious beams Diaphragms Brewster B1 B5 Brewster link axis Distance from BS (m) Distance from west axis (cm) One spurious beam

48 48 With the new brewster One of the predicted spurious beams Diaphragms Brewster B1 B5 Brewster link axis Distance from BS (m) Distance from west axis (cm) One spurious beam

49 49 With the new brewster Considering beam diameter defined as 2 w 0 One of the predicted spurious beams Diaphragms Brewster B1 B5 Brewster link axis Distance from BS (m) Distance from west axis (cm) One spurious beam

50 50 Additional diaphragm before the Brewster Added from SR tower A possible mitigation Diaphragms Brewster B1 B5 Brewster link axis Distance from BS (m) Distance from west axis (cm) One spurious beam

51 51 Scattering of spurious light from the ITF ? Yes, but …. …… no difference in the ITF noise

52 52 Damping the Brewster

53 53 Injection with shaker: comparison with old Brewster Noise injections (May 9 th )

54 54 Injection with loudspeaker: comparison with old Brewster Noise injections (May 9 th )

55 55 Tapping tests on injection Brewster, injection tower, injection table - Small effects on ITF sensitivity - No excitation of ITF noise structures/resonances - Much less sensitive than detection tower and Brewster Tapping tests on injection Brewster, injection tower, injection table - Very small effects on ITF sensitivity - No excitation of ITF noise structures/resonances Tapping test attempted on NI - failed due to unlock Need to study sensitivity to all towers vibration - need to learn how to excite towers without unlock - place seismometers in more useful positions (now measuring vertical seismic noise near tower base) More investigations (IV)

56 56 “Mystery noise”

57 57 “Mystery noise”

58 58 “Mystery noise”

59 59 “Mystery noise”

Download ppt "1 Virgo commissioning progress Post VSR1 plans E. Tournefier STAC meeting June 26 th,2007."

Similar presentations

Ads by Google